The winner of the first annual Tom Morris Golf Invitational won $160 in the competition which was held in 1910. In 2015, the winner received $1,570,000. If the winner's purse continues to increase at the same interest rate, how much will the winner receive in 2044?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

1,570,000=160*(1+r/100)^105

(1,570,000/160)^(1/105)=(1+r/100)

(1+r/100)=1.0915

r=1.0915-1

=0.0915

A=$1,570,000*(1+0.0915)^29

=$19878846.50(Approx).